OpenStudy (anonymous):

sin theta - tan theta cos theta + cos (pi/2 - theta)

OpenStudy (calculusfunctions):

Is your question, to simplify the above expression?

OpenStudy (calculusfunctions):

If so then you must apply the identities.

OpenStudy (anonymous):

cot ( theta - pi / 2 )= cos ( theta - pi / 2 ) / sin ( theta - pi / 2 ) cos ( theta - pi / 2 )= cos(theta) cos (pi/2) + sin(theta) sin (pi/2) = cos(theta) (0) + sin(theta) (1) = sin(theta) sin ( theta - pi / 2 )= sin(theta) cos(pi/2) - cos(theta) sin(pi/2) = sin(theta) (0) - cos(theta) (1) = -cos(theta) so you get, cot ( theta - pi / 2 )= sin(theta) / -cos(theta) cot ( theta - pi / 2 )= -tan(theta)

OpenStudy (anonymous):

i hope that is right :-)

OpenStudy (anonymous):

u guys can also recheck

OpenStudy (calculusfunctions):

\[\sin (\frac{ \pi }{ 2 }-\theta)=\cos \theta \] \[\cos(\frac{ \pi }{ 2 }-\theta)=\sin \theta\] \[\tan (\frac{ \pi }{ 2 }-\theta)=\cot \theta \]etc.